Exercise 1-4 Fixed and Variable Cost Behavior [L01-4] Espresso Express operates a number of espresso coffee stands in busy suburban malls. The fixed weekly expense of a coffee stand is $1,900 and the variable cost per cup of coffee served is $0.22 1. Fill Iin the following table with your estimates of the companys total cost and average cost per cup of coffee at the indicated levels of activity. 2. Does the average cost per cup of coffee served increase, decrease, or remain the same as the number of cups of coffee served in a week increases? Complete this question by entering your answers in the tabs below. Required 1 Required 2 Fill in the following table with your estimates of the companys total cost and average cost per cup of coffee at the indicated levels of activity. (Round the Average cost per cup of coffee served to 3 decimal places.) Cups of Coffee Served in a Week 900 1,000 1,1 Fixed cost Variable cost Total cost Average cost per oup of coffee served Required 2>
Answer #1
cups of coffee served in a week
900 1000 1100
Fixed cost 1900 1900 1900
Variable cost 198    [900*.22] 220    [1000*.22] 242     [1100*.22]
Total cost 2098 2120 2142
Average cost per cup of coffee served 2098/900=2.331 2120/1000=2.120 2142/1100=$ 1.947

2)

Average cost per cup of coffee has decreased with increase in number of cups served.